OpenStudy (anonymous):

What is a possible first step to eliminate a variable in the following system? 7x – 9y = 2 3x – 6y = –4 A. Multiply the first equation by 2 and the second equation by –3. B. Multiply the first equation by –3 and the second equation by –2. C. Multiply the first equation by 7 and the second equation by 3. D. Multiply the first equation by –2 and the second equation by 4.

OpenStudy (anonymous):

7x-9y=2 3x-6y=-4 When doing elimination the idea is to get one of the variables to cancel out. 2(7x-9y)=2(2) = 14x-18y=4 [first equation] -3(3x-6y)=-3(-4) = -9x+18y=-12 [second equation] You'll notice that the y's in this case cancel each other out. Multiply the first equation by 2 and the second equation by -3

OpenStudy (anonymous):

ThankYou!! :)